# LOLA: LLM-Assisted Online Learning Algorithm for Content Experiments

### v2

@@ Abstract
In the rapidly evolving digital content landscape, media firms and news publishers require automated and efficient methods to enhance user engagement. This paper introduces the LLM-Assisted Online Learning Algorithm (LOLA), a novel framework that integrates Large Language Models (LLMs) with adaptive experimentation to optimize content delivery. Leveraging a large-scale dataset from Upworthy, which includes 17,681 headline A/B tests, we first investigate three pure-LLM approaches: prompt-based methods, embedding-based classification models, and fine-tuned open-source LLMs. We find that prompt-based approaches perform poorly, achieving no more than 65% accuracy in identifying the catchier headline. In contrast, both OpenAI-embedding-based classification models and fine-tuned Llama-3 with 8 billion parameters achieve an accuracy of around 82-84%. We then introduce LOLA, which combines the best pure-LLM approach with the Upper Confidence Bound algorithm to allocate traffic and maximize clicks adaptively. Our numerical experiments on Upworthy data show that LOLA outperforms the standard A/B test method (the current status quo at Upworthy), pure bandit algorithms, and pure-LLM approaches, particularly in scenarios with limited experimental traffic. Our approach is scalable and applicable to content experiments across various settings where firms seek to optimize user engagement, including digital advertising and social media recommendations.
@@ 1 Introduction
To address this question, we leverage a large dataset of content experiments by the publishing firm Upworthy conducted in the 2013–2015 timeframe. The dataset consists of 17,618 headline tests spanning 77,245 headlines, 277 million impressions, and 3.7 million clicks. The goal of each A/B test was to identify the best headline among a set of candidate headlines (for a given article). This dataset is an excellent test-bed for our study for a few reasons – (1) it is based on an extremely large number of A/B tests over a high volume of impressions, (2) each A/B test received a relatively large number of impressions, allowing for accurate estimates of each headline’s performance, and (3) it is sourced from a real media firm and represents the actions of a very large number of readers over a sufficiently long period to offer meaningful insights.
In the first part of the paper, we use this dataset to examine the extent to which pure-LLM-based approaches can accurately predict headline attractiveness. For this exercise, we focus on pairs of headlines (within a given A/B test) whose CTRs are significantly different (based on the observed data). This gives us a total of 39,158 pairs of headlines and allows us to recast the prediction task as a binary classification task, where a random guess gives 50% accuracy, and the perfect classifier reaches 100% accuracy. Thus, this serves as a preliminary testbed to evaluate whether pure-LLM-based approaches can predict which of the two headlines is catchier (or clickable).
@@
We consider the three most widely used LLM-based approaches for this analysis—(1) Prompt-based approaches, (2) LLM text-embedding approaches, and (3) Fine-tuning approaches. For prompt-based methods, we consider two approaches: (a) Zero-shot prompting and (b) In-context learning. Zero-shot prompting is a technique in which an LLM generates responses without being explicitly trained on specific examples. This is the most common way in which most users interact with the LLM. In-context learning involves providing the LLM with a few demonstrations of similar tasks before generating responses for a specific task (min2022rethinking; xie2021explanation). We find that GPT-3.5 performs as poorly as a random guess for both prompt-based approaches, with accuracy hovering around 50%. GPT-4 performs marginally better—it achieves an accuracy of 57.26% for zero-shot prompting and 60-65% for in-context learning. Overall, this suggests that prompt-based approaches cannot aid/replace content experiments in any meaningful way.
@@
Next, we turn to the second approach: text embeddings. Specifically, we transform headlines into embedding vectors using the most recent embedding model from OpenAI. We then concatenate two embeddings as inputs, along with the winner with higher CTR as the output, to train binary classification models, including logistic regression and multilayer perceptron (a type of neural network). We find that OpenAI’s embeddings combined with simple logistic regression can achieve around 82.50% accuracy, while more complex neural networks do not improve performance over the simple logistic model. Finally, we fine-tune a state-of-the-art open-source LLM — Llama-3 with 8 billion parameters, using low-rank adaptation (LoRA) (hu2021lora). This approach performs the best, with 83.40% accuracy, which is a small improvement over the embedding-based approach. However, even the best-fine-tuned LLMs are unable to perfectly generate/predict which types of content are most appealing to users. As such, none of the LLM-based approaches can match the accuracy of experimentation-based approaches.
@@
Therefore, in the second part of the paper, we propose and evaluate a novel framework for content experiments that combines the strengths of LLM-based approaches with the advantages of online learning algorithms, termed LOLA (LLM-Assisted Online Learning Algorithm). Our key insight is to use predictions from an LLM model as priors in online algorithms to optimize experimentation. The approach involves two steps. First, we train an LLM-based CTR prediction model using LoRA fine-tuning. The second step integrates these predictions into online learning algorithms. LOLA is a general framework, and depending on the specific goals of the experimentation, it can be adapted for different prediction models in the first step and different online algorithms in the second step. For example, if the goal is to minimize regret (i.e., maximize total reward/clicks), we propose a modified Upper Confidence Bound algorithm called LLM-Assisted 2-Upper Confidence Bounds (LLM-2UCBs) that is detailed in §$\S$4.2. The idea of LLM-2UCBs is intuitive: we can view LLM-based CTR predictions as auxiliary samples before the start of the online algorithm (e.g., a 1% CTR can be viewed as 100 impressions with 1 click or 1,000 impressions with 10 clicks). After fine-tuning this auxiliary impression size hyperparameter, we incorporate these auxiliary samples into the standard UCB algorithm (auer2002finite) to obtain the reward estimator and shrink the upper confidence bound.
@@
In summary, our paper makes two key contributions to the literature. First, from a methodological perspective, we develop a novel framework, LOLA, to leverage LLMs for improving content experimentation in digital settings. To the best of our knowledge, this is the first paper that proposes combining LLMs with adaptive experimentation techniques. We demonstrate the value of our approach using a large-scale dataset from Upworthy when compared to baseline methods. Second, from a managerial perspective, LOLA can be used in a broad range of settings where firms need to decide which content to show. While we focus on the news/publishing industry given our empirical context, the LOLA framework is general and can easily extend to other settings, including digital advertising, email marketing, and website design. Further, the method is cost-effective, can be built on open-source LLMs (alleviating data privacy concerns), and is easily deployable across different domains once fine-tuned on relevant datasets.
@@ 1 Why Yoko Ono Is The Only Thing Standing Between New York And Catastrophic Gas Fracking 2,734 34
# of headline pairs 140,655
# of headline pairs with different CTRs at 0.05 significant level 39,158
# of headline pairs with different CTRs at 0.01 significant level 23,682
@@
To examine whether these A/B tests were informative, or not, we generate pairs of headlines and conduct pairwise t-tests on the CTRs of the headlines. For example, if four headlines were tested for one article, we generate $4\times 3/2=6$ pairs of headlines, and test whether each pair is significantly different from each other; [^3] The issue of multiple testing may arise when conducting a large number of pairwise comparisons, as it increases the likelihood of Type I errors, or false positives. In our dataset, tests involving 10 pairs or less (equivalent to no more than 5 packages) account for over 80% of all tests. The risk of multiple testing is relatively low when conducting fewer than 10 tests, and conventional significance levels, e.g., $\alpha=0.05$ are generally adequate to account for this risk. Ultimately, our objective is to develop a more effective decision rule for assigning experimental traffic, and in §$\S$4, we utilize all available data without concern for multiple testing. see Table 4 for examples of such tests for the two A/B tests highlighted earlier in Table 1. After generating pairs, we obtain 140,655 comparable headline pairs across all tests, of which 39,158 pairs significantly differ in CTRs at a 0.05 level and 23,682 at a 0.01 level. In summary, we find that only 28% of the tests were significant at the 0.05 level, and a mere 17% are significant at the 0.01 level. Overall, these results suggest that most pairs of headlines tend to perform similarly, and the task of learning which headline is catchier is non-trivial. To further illustrate the difficulty of this task, we conduct a survey, where we give human users pairs of headlines from the set of headline pairs that are significantly different from each other and ask them to identify the catchier headline. Even within this set of headline pairs, we find that the respondents’ accuracy in the survey was not significantly better than random guessing. This further highlights the fact that the task of identifying engaging content is inherently challenging, even for humans. See Web Appendix §$\S$§A for more details of the survey and its results.
Test ID Headline 1 Headline 2 CTR 1 CTR 2 p value Winner
1 How YOU Can Help… Why Yoko Ono Is… 0.0071 0.0124 0.054 2
1 Why Yoko Ono Is… New York’s Last… 0.0124 0.0056 0.009 1
1 How YOU Can Help… New York’s Last… 0.0071 0.0056 0.496 1
2 If You Know Anyone Who… Hey Dude. If… 0.0289 0.010 0.000 1
Table 4: Example of pairwise comparisons for two A/B tests from the Upworthy data.
@@ 3 Pure-LLM-Based Methods
In this section, we consider the 39,158 headline pairs that significantly differ from each other at the 0.05 significance level. We focus on these significant pairs for three reasons: First, insignificant pairs suggest that decisions in final headline selection do not substantially affect users’ click behavior, implying that accuracy in predicting outcomes for “equally good” samples is not critically important from a business perspective. Second, the primary objective here is to evaluate various LLM-based classification methods to determine the most effective approach for the algorithm design in §$\S$4. As such, relative comparison across methods is more important as long as we use the same dataset for all methods. Third, as demonstrated in Web Appendix §$\S$§D.2.1, our robustness checks confirm that including insignificant pairs in the training phase does not affect the overall performance of the methods considered. Furthermore, we note that all methods discussed in this section can be readily adapted from binary to multi-class classification tasks. Indeed, in §$\S$4 where we introduce our LOLA approach, we incorporate practical elements such as multiple headlines and insignificant pairs to ensure our experimental setup closely mirrors real-world conditions and that our conclusions are directly applicable to business scenarios.
In the rest of this section, we investigate the performance of three widely used LLM-based approaches—(1) GPT prompt-based approaches, including zero-shot prompting and in-context learning, (2) embedding-based classification models, and (3) fine-tuning of LLMs.
@@ 3.1 Prompt-based Approaches
Prompt-based approaches use GPT prompts to learn which headline is catchier, and there is no modeling or explicit fine-tuning involved. Prompt-based approaches have been the standard use-case for AI models in the marketing literature so far and have been used to examine if and when LLMs can simulate consumer behavior in market research studies (gui2023challenge; brand2023using; li2024frontiers). We consider two prompt-based approaches: (1) Zero-shot prompting and (2) In-context learning.
@@ 3.1.1 Zero-shot Prompting
Zero-shot prompting is a technique in which an LLM generates responses or performs tasks without being explicitly trained on specific examples. We evaluate the performance of headline pair comparison using zero-shot prompting based on OpenAI’s GPT models. We do this through OpenAI’s API in Python, which offers various parameters to customize and control the model’s behavior and responses. The key parameters include specifying which GPT model to use (such as GPT-3 or GPT-4) and the system parameters, which guide the type of responses generated. The user’s input or query is contained in the user parameter, while the assistant parameter holds the model’s previous responses to maintain context in an ongoing conversation. Additionally, we set the default temperature parameter of 1.444The temperature parameter scales and controls the randomness of the model’s responses. A larger value results in a more diverse and creative output, while a lower temperature makes the output more deterministic and focused. For simplicity, we use the default setting with temperature as 1.0. We also tested a lower temperature of 0.1 to generate more deterministic outcomes, but we did not observe any significant difference in accuracy.
@@
We design the zero-shot prompting to ask GPT which headline is catchier, as shown in Figure 1. We test this prompt using three OpenAI large language models, [^5] As recommended by OpenAI, see this page, accessed on May 24, 2024, one can default to GPT-3.5-turbo, GPT-4-turbo, or GPT-4o. OpenAI notes that GPT-4-turbo and GPT-4o offer similar levels of intelligence. each differing in parameter size: (1) GPT-3.5 (GPT-3.5-turbo-0125), which uses around 175 billion parameters, (2) GPT-4 (GPT-4-turbo-2024-04-09) which uses approximately 1.7 trillion parameters, and (3) GPT-4o (GPT-4o-2024-05-13), which uses an even larger number of parameters. The goal here is to investigate whether improvements in model architecture and parameter size can improve the accuracy of our predictions via standard prompts.
@@ 3.1.2 In-context Learning
We design in-context learning prompts as shown in Figure 2. Our prompts include demonstrations of catchier headlines and ask the model to determine which headline is catchier for a subsequent pair of headlines. We use the headline with the higher CTR from the test, as indicated in the “Winner” column of Table 4, as the correct answer in our demonstrations. A few demonstrations are typically sufficient to guide the model. Improvement tends to have diminishing returns quickly, i.e., adding more demonstrations does not significantly enhance performance after a certain point (min2022rethinking). In our numerical experiments, we use two or five demonstrations to investigate the impact of the number of examples on performance. [^6] We did run experiments with a larger number of demonstrations, but did not find any significant improvements from doing so; and hence do not report them here.
An interesting aspect of in-context learning is the model’s robustness to incorrect labels in the demonstrations. (min2022rethinking) show that randomized labels in demonstrations barely affect performance and highlight other critical aspects like the label space, input text distribution, and sequence format. This is likely because the model relies more on the distribution of input text, label space, and format rather than specific input-label mappings. That is, in-context learning seems to benefit from recognizing general patterns and structures in the data, which the model infers from its pre-training rather than the actual labels. To examine if this is the case in our setting as well, we also consider experiments where we intentionally flip the labels in the demonstrations provided to the model.

### v3

@@ Abstract
Modern media firms require automated and efficient methods to identify content that is most engaging and appealing to users. Leveraging a large-scale dataset from Upworthy (a news publisher), which includes 17,681 headline A/B tests, we first investigate the ability of three pure-LLM approaches to identify the catchiest headline: prompt-based methods, embedding-based methods, and fine-tuned open-source LLMs. Prompt-based approaches perform poorly, while both OpenAI-embedding-based models and the fine-tuned Llama-3-8B achieve marginally higher accuracy than random predictions. In sum, none of the pure-LLM-based methods can predict the best-performing headline with high accuracy. We then introduce the LLM-Assisted Online Learning Algorithm (LOLA), a novel framework that integrates Large Language Models (LLMs) with adaptive experimentation to optimize content delivery. LOLA combines the best pure-LLM approach with the Upper Confidence Bound algorithm to allocate traffic and maximize clicks adaptively. Our numerical experiments on Upworthy data show that LOLA outperforms the standard A/B test method (the current status quo at Upworthy), pure bandit algorithms, and pure-LLM approaches, particularly in scenarios with limited experimental traffic. Our approach is scalable and applicable to content experiments across various settings where firms seek to optimize user engagement, including digital advertising and social media recommendations.
@@ 1 Introduction
To address this question, we employ a large dataset of content experiments by the publishing firm Upworthy conducted in the 2013–2015 timeframe. The dataset consists of 17,618 headline tests spanning 77,245 headlines, 277 million impressions, and 3.7 million clicks. The goal of each A/B test was to identify the best headline among a set of candidate headlines (for a given article). This dataset is an excellent test-bed for our study for a few reasons – (1) it is based on an extremely large number of A/B tests over a high volume of impressions, (2) each A/B test received a relatively large number of impressions, allowing for accurate estimates of each headline’s performance, and (3) it is sourced from a real media firm and represents the actions of a very large number of readers over a sufficiently long period to offer meaningful insights.
In the first part of the paper, we use this dataset to examine the extent to which pure-LLM-based approaches can accurately predict headline attractiveness. For this exercise, we test the different approaches’ accuracies, defined as the success rate of identifying the headline with the highest click-through rate (CTR) in a given A/B test.
@@
We consider the three most widely used LLM-based approaches for this analysis—(1) Prompt-based approaches, (2) LLM text-embedding approaches, and (3) Fine-tuning approaches. For prompt-based methods, we consider two approaches: (a) Zero-shot prompting and (b) In-context learning. Zero-shot prompting is a technique in which an LLM generates responses without being explicitly trained on specific examples. This is the most common way in which most users interact with the LLM. In-context learning involves providing the LLM with a few demonstrations of similar tasks before generating responses for a specific task (min2022rethinking; xie2021explanation). We find that GPT-3.5 performs as poorly as a random guess for both prompt-based approaches. GPT-4 performs marginally better—it achieves an accuracy of 37.85% for zero-shot prompting and 38-40% for in-context learning. Overall, this suggests that prompt-based approaches cannot aid/replace content experiments in any meaningful way.
@@
Next, we turn to the second approach: text embeddings. Specifically, we transform headlines into embedding vectors using the most recent embedding model from OpenAI. We then use the embedding as the input, along with the CTR as the output, to train CTR prediction models, including linear regression and multilayer perceptron (a type of neural network). After getting predicted CTRs, we choose the headline with the highest predicted CTR as the winner. We find that OpenAI’s embeddings combined with simple linear regression can achieve around 46.28% accuracy, while more complex neural networks do not improve performance over the simple linear model. Finally, we fine-tune a state-of-the-art open-source LLM — Llama-3 with 8 billion parameters, using low-rank adaptation (LoRA) (hu2021lora). This approach performs the best, with 46.86% accuracy, which is a small improvement over the embedding-based approach. However, even the best-fine-tuned LLMs are unable to perfectly predict which types of content are most appealing to users. As such, none of the LLM-based approaches can match the accuracy of experimentation-based approaches.
Therefore, in the second part of the paper, we propose and evaluate a novel framework for content experiments that combines the strengths of LLM-based approaches with the advantages of online learning algorithms, termed LOLA (LLM-Assisted Online Learning Algorithm). Our main insight is to use predictions from an LLM model as priors in online algorithms to optimize experimentation. While the general idea that priors can impact regret and rewards has existed for some time (bubeck2013prior; russo2014learning), these papers typically assume that the quality and distribution of priors are known. However, a key challenge in our setting is that the LLM predictions cannot be naively treated as priors since we do not know their prediction error or theoretical properties. Therefore, we build on the recently proposed 2-UCBs algorithm by gur2022adaptive to overcome this challenge. The original 2-UCBs algorithm was designed for a setting where the experimenter has access to auxiliary data and knows both the size of this auxiliary sample ($n^{aux}$) and knowledge of the prediction error in this sample. Our innovation is to extend this algorithm to accommodate LLM predictions by treating these predictions as if they came from a “pseudo-sample” and then to fine-tune the size of this pseudo-sample for a given application based on prior data using hyper-parameter tuning. Thus, we are able to combine LLM predictions with online experimentation without making strong assumptions on the theoretical properties of the LLM-based predictors.
@@
LOLA involves two steps. First, we train an LLM-based CTR prediction model using LoRA fine-tuning. The second step integrates these predictions into online learning algorithms. LOLA is a general framework, and depending on the specific goals of the experimentation, it can be adapted for different prediction models in the first step and different online algorithms in the second step. For example, if the goal is to minimize regret (i.e., maximize total reward/clicks), we propose a modified Upper Confidence Bound algorithm called LLM-Assisted 2-Upper Confidence Bounds (LLM-2UCBs) that is detailed in §$\S$4.2. The idea of LLM-2UCBs is intuitive: we can view LLM-based CTR predictions as auxiliary samples before the start of the online algorithm (e.g., a 1% CTR can be viewed as 100 impressions with 1 click or 1,000 impressions with 10 clicks). After fine-tuning this auxiliary impression size hyperparameter, we incorporate these auxiliary samples into the standard UCB algorithm (auer2002finite) to obtain the reward estimator and shrink the upper confidence bound.
@@
In summary, our paper makes three key contributions to the literature. First, from an empirical perspective, the paper provides evidence that LLMs (or, more precisely, the current vintage of LLMs) cannot replace experimental approaches in the media and digital marketing industry. Second, from a methodological perspective, we develop a novel framework, LOLA, to leverage LLMs for improving content experimentation in digital settings. To the best of our knowledge, this is the first paper that proposes combining LLMs with adaptive experimentation techniques. We demonstrate the value of our approach using a large-scale dataset from Upworthy when compared to baseline methods. Our work thus provides a foundation for future research into more advanced algorithms and offers insights for real-world validation in industry settings. Third, from a managerial perspective, LOLA can be used in a broad range of settings where firms need to decide which content to show. While we focus on the news/publishing industry, given our empirical context, the LOLA framework is general and can easily extend to other settings, including digital advertising, email marketing, and website design. Further, the method is cost-effective, can be built on open-source LLMs (alleviating data privacy concerns), and is easily deployable across different domains once fine-tuned on relevant datasets.
@@ 1 Why Yoko Ono Is The Only Thing Standing Between New York And Catastrophic Gas Fracking 2,734 34
To illustrate that learning which headline is best is a non-trivial task, we conduct a survey where we give human users pairs of headlines from the set of headline pairs that are significantly different from each other and ask them to identify the catchier headline. Even within this set of headline pairs, we find that the respondents’ accuracy in the survey was not significantly better than random guessing. This further highlights the fact that the task of identifying engaging content is inherently challenging, even for humans. See Web Appendix §$\S$§A for more details of the survey and its results. [^3] Note that just because an average reader cannot discern/identify the most engaging or catchiest headline among a pair/set of headlines, it does not follow that such differences do not exist. Indeed, the whole premise of the experimentation culture in digital firms relies on the idea that while managers/lay users cannot correctly predict which version of an ad creative/promotion/website layout will be more engaging, such differences exist and that identifying these differences and implementing the best-performing arm is important and will lead to better consumer engagement and business outcomes.
@@ 3 Pure-LLM-Based Methods
We divided the 17,681 headline tests into three folds: 70% for training, 10% for hyperparameter fine-tuning of the algorithms (as discussed in §$\S$4), and 20% for testing. This sample split is used consistently throughout the remainder of the paper unless stated otherwise. Specifically, in this section, we have 12,376 headline tests as a training set and 3,263 headline tests as a test set. [^4] Note this test set is slightly smaller than 20% of headline tests. This is because, to avoid information leakage, we deleted the duplicated tests that had the same headlines in the training dataset. This happens because Upworthy allowed additional experiments for the same headline under a new experiment ID, and the editor may require an extra test to make an informed decision.
In the rest of this section, we investigate the performance of three widely used LLM-based approaches—(1) GPT prompt-based approaches, including zero-shot prompting and in-context learning, (2) embedding-based models, and (3) fine-tuning of LLMs.
@@ 3.1 Prompt-based Approaches
Prompt-based approaches use GPT prompts to select which headline is the catchiest, and there is no modeling or explicit fine-tuning involved. Prompt-based approaches have been the standard use-case for AI models in the marketing literature so far and have been used to examine if and when LLMs can simulate consumer behavior in market research studies (gui2023challenge; brand2023using; li2024frontiers). We consider two prompt-based approaches: (1) Zero-shot prompting and (2) In-context learning.
@@ 3.1.1 Zero-shot Prompting
Zero-shot prompting is a technique in which an LLM generates responses or performs tasks without being explicitly provided by any specific examples. We evaluate the performance using zero-shot prompting based on OpenAI’s GPT models. We do this through OpenAI’s API in Python, which offers various parameters to customize and control the model’s behavior and responses. The key parameters include specifying which GPT model to use (such as GPT-3 or GPT-4) and the system parameters, which guide the type of responses generated. The user’s input or query is contained in the user parameter, while the assistant parameter holds the model’s previous responses to maintain context in an ongoing conversation. Additionally, we set the temperature parameter to be 0, which indicates there is no randomness in the GPT’s response for reproducibility. [^5] The temperature parameter scales and controls the randomness of the model’s responses. A larger value results in a more diverse and creative output, while a lower temperature makes the output more deterministic and focused. The default setting with temperature is 1.0. We also tested the default temperature, and we did not observe any significant difference in accuracy.
@@
We design the zero-shot prompting to ask GPT which headline is the catchiest, as shown in Figure 1. We tell ChatGPT that it is an expert in marketing and specializing in optimizing headline selection. In this figure, we use letters (A,B,C,...) to mark different headlines. (brucks2023prompt) shows that different marker types and the order of options provided to LLM might affect the model’s response, in the sense that LLM might prefer to select the first option or select a certain marker. To mitigate this issue, we also consider other markers, numbers (1,2,3,...) and symbols (!,@,#,...), and we randomize the order of the options. Further, we adopt the common practice of LLM prompting for a specialized task by detailing the task for the LLM and clearly defining the meaning of a catchy headline. Notably, the first paragraph of the prompt assigns the role of a digital media expert to the LLM. This technique, commonly referred to as role prompting, involves explicitly assigning a specific role to the LLM to guide its responses. Role prompting has been shown to be an effective strategy in the broader survey of prompt engineering techniques (schulhoff2024prompt), and we find that it works well in our setting, too.
We test this prompt using three OpenAI large language models, [^6] As of May 24, 2024, OpenAI recommends defaulting to GPT-3.5-turbo, GPT-4-turbo, or GPT-4o. OpenAI notes that GPT-4-turbo and GPT-4o offer similar levels of intelligence. each differing in parameter size: (1) GPT-3.5 (GPT-3.5-turbo-0125), which uses around 175 billion parameters, (2) GPT-4 (GPT-4-turbo-2024-04-09) which uses approximately 1.7 trillion parameters, and (3) GPT-4o (GPT-4o-2024-05-13), which uses an even larger number of parameters. The goal here is to investigate whether improvements in model architecture and parameter size can improve the accuracy of our predictions via standard prompts.
@@ 3.1.2 In-context Learning
We design in-context learning prompts as shown in Figure 2. Our prompts include demonstrations of best-performing headlines and ask the model to evaluate the subsequent headlines. The demonstrations are selected from the training set, and for different prompt inquiries, we use the same demonstration to maintain comparability. A few demonstrations are typically sufficient to guide the model. Improvement tends to have diminishing returns quickly, i.e., adding more demonstrations does not significantly enhance performance after a certain point (min2022rethinking). In our numerical experiments, we use two or five demonstrations to investigate the impact of the number of examples on performance. [^7] We did run experiments with a larger number of demonstrations, but did not find any significant improvements from doing so; and hence do not report them here.
An interesting aspect of in-context learning is the model’s robustness to incorrect labels in the demonstrations. (min2022rethinking) show that providing randomly selected labels as the answers in demonstrations barely affects performance and highlight other critical aspects like the label space, input text distribution, and sequence format. This is likely because the model relies more on the distribution of input text, label space, and format rather than specific input-label mappings. That is, in-context learning seems to benefit from recognizing general patterns and structures in the data, which the model infers from its pre-training rather than the actual labels. To examine if this is the case in our setting as well, we also consider experiments where we intentionally provide incorrect labels in the demonstrations.